HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025  Not logged in ELOG logo
Message ID: 69     Entry time: Wed Apr 2 07:41:10 2025
Author: Michael, Alex 
Category: General 
Subject: Frequency shift in Schottky 
After the beam  was restored at 07:25, we think that the frequency peak in the 244 MHz RSA2 has jumped again (closer to the center). 
Previous to that, during the longer break in the previous night, the peak jumped to higher frequencies.
ELOG V3.1.5-fc6679b